Synthetic Division:

  1. (3.) (3x^5 - 2x^3 + 5x^2 - x+3) / (3x-2)
1
Expert's answer
2021-09-20T05:03:10-0400

"\frac{(3x^5 - 2x^3 + 5x^2 - x+3) }{ (3x-2)}"





"=x^4+\frac{2x^3}{3}-
\frac{2x^2}{9}+
\frac{41x}{27}+
\frac{55}{81}+
\frac{\frac{353}{81}}{3x-2}"